How to solve by elimination: 5x-7y=-28 2x-2y=-12

How to solve by elimination 9x-9y=0 6x-2y=-20

Answer by TimothyLamb(4379)

5x-7y=-28
2x-2y=-12
---
put the system of linear equations into standard form
---
5x-7y=-28
2x-2y=-12
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = -7
y = -1
---
by elimination:
(5x-7y=-28)*(2)
(2x-2y=-12)*(-7)
add the resulting equations, eliminating the y term.
solve for x
plug x into either equation and solve for y
---
---
9x-9y=0
6x-2y=-20
---
put the system of linear equations into standard form
---
9x-9y=0
6x-2y=-20
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = -5
y = -5
---
by elimination:
(9x-9y=0)*(2)
(6x-2y=-20)*(-9)
add the resulting equations, eliminating the y term.
solve for x
plug x into either equation and solve for y
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php